OpenStudy (anonymous):

x^3-6x^2+25x in complex?

OpenStudy (anonymous):

and x^3+9x^2+22x

OpenStudy (anonymous):

can you factor out anything here?

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

correct, also if you are solving for the roots of an EQUATION i suggest you augment =0 to your equations.

OpenStudy (anonymous):

but it wants i to be in the factored form, in complex, how do i do that

OpenStudy (anonymous):

what do you get if you factor out a zero?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

lol i meant factor out an x

OpenStudy (anonymous):

whats your equation?

OpenStudy (anonymous):

x^3-6x^2+25x which would be x(x^2-6x+25)

OpenStudy (anonymous):

now factor the x^2-6x+25 part

OpenStudy (anonymous):

uhm how

OpenStudy (anonymous):

you could use the quadratic equation? \[\frac{-b \pm \sqrt{b^2 - (4ac)}}{2a}\]

OpenStudy (anonymous):

i got (6 plusorminus √-64) / 2 so how do i put that in parenthesis

OpenStudy (anonymous):

\[\frac{6}{2} \pm \frac{\sqrt{-64}}{2}\] can you simplify this?

OpenStudy (anonymous):

3plusorminus √-64/2?

OpenStudy (anonymous):

yes now can you simplify \[\frac{\sqrt{(-64)}}{2} \] what is -64 ? 64 times what?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

-i?

OpenStudy (anonymous):

-1 is correct so it looks like this: \[\sqrt{(64)\times(-1)}\] and you were asking about complex numbers well do you know the definition of i and can you apply it here?

OpenStudy (anonymous):

this is the same as \[\sqrt{64} \times \sqrt{-1} \]

OpenStudy (anonymous):

so how would it be?

OpenStudy (anonymous):

well what is the square root of 64?

OpenStudy (anonymous):

oh 8

OpenStudy (anonymous):

also what is the square root of -1?

OpenStudy (anonymous):

and whats \[\frac{8}{2}\]

OpenStudy (anonymous):

4

OpenStudy (anonymous):

ok write out what you have reduced it to thus far?

OpenStudy (anonymous):

3plusorminus√-8

OpenStudy (anonymous):

\[3\pm \sqrt{-1}\times 8\] this is what you got, so how to we take care of the \[\sqrt{-1}\]

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

google it, you will be surprised!

OpenStudy (anonymous):

i?

OpenStudy (anonymous):

w00t YES ISN'T THAT EXCITING!

OpenStudy (anonymous):

so whats the final solution?

OpenStudy (anonymous):

3plusorminusi√8?

OpenStudy (anonymous):

why do you keep putting \[\sqrt{8}\] in there?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

we said that \[\sqrt{64}\] was 8 right? So therefore its just 8 but then its \[\frac{8}{2}\] was it not?

OpenStudy (anonymous):

so 8i?

OpenStudy (anonymous):

check again its not 8, write out the entire simplified solution for me pelase!

OpenStudy (anonymous):

--_____---

OpenStudy (anonymous):

you are almost there dont give up go back and look at the conversation!!!!!

OpenStudy (anonymous):

i fel like giving up

OpenStudy (anonymous):

well dont you are almost there!!!!! keep trying!

OpenStudy (anonymous):

can you please just lay it out for me?